# Andreas Petrou-Zeniou

### v3

@@ 3 Inference on Multiple Winners in the JOBSTART Demonstration
In this section, we revisit the JOBSTART demonstration due to Cave et al. (1993) and subsequent replication failure in Miller et al. (2005), which has been previously studied in the selective inference literature by Andrews et al. (2023). The JOBSTART demonstration was a randomized controlled trial taking place between 1985 and 1988 in 13 sites, with the intention of studying the effects of a vocational training program on the employment outcomes of young, low-skilled school dropouts. The treatment group was given access to a suite of JOBSTART services which were inaccessible to those in the control group. Among the sites included in the JOBSTART study, only one site, the Center for Employment Training (CET) in San Jose, saw a large and statistically significant estimate of the effect on earnings. Cave et al. (1993) note that they cannot attribute the unique success of CET to a particular feature of the program, but suggest that the CET’s strong connections with San Jose employers, or their robust placement efforts, may explain some of the value-add of the program. Motivated by the success of the CET program, Miller et al. (2005) replicate the intervention at 12 sites. They find that, even in replication sites deemed to have high fidelity to the original CET program of Cave et al. (1993), the estimated effect of the program’s services on enrollees’ earnings was not statistically significant.
@@
Andrews et al. (2023) study the possibility that this replication failure is due to a winner’s curse. In particular, they consider the possibility that the estimates of the effect of CET on earnings in the original JOBSTART demonstration are upwardly biased by virtue of CET being the site with the largest estimated effect. We consider a complementary thought experiment. In particular, since the exact mechanism by which Miller et al. (2005) select a program for replication is unknown, we consider the possibility that the replication failure between the two studies can be explained by an alternative selection rule. In particular, we consider the possibility that the sites selected by Miller et al. (2005) are chosen according to a statistical significance cutoff. Under this selection rule, it is impossible for the econometrician to know ex-ante how many sites will be selected for replication. When multiple selections are made, the conditional and hybrid approaches of Andrews et al. (2022) are difficult to apply, as we discuss in section 4.1. When only one selection is made, one must consider an alternative selection rule to that considered by Andrews et al. (2023). As a result, we suggest that our methods are better-suited for inference in this particular setting.
@@
Empirically, we show that our two-step confidence regions for the effect of the CET program on earnings in the Cave et al. (1993) study are wider under this alternative selection rule than under the selection rule considered by Andrews et al. (2023). Nevertheless, as in Andrews et al. (2023), these confidence regions still exclude zero. This finding suggests that a winner’s curse does not fully explain the replication failure between the studies of Cave et al. (1993) and Miller et al. (2005). Our two-step confidence regions provide more informative inference than projection confidence regions in this application, To be precise, our two-step confidence regions are between 6% and 8% shorter than the corresponding projection confidence regions. Hybrid and conditional inference tend to provide narrower confidence sets than the two-step approach, though only in cases where a single site is selected for replication. Finally, we study the frequency with which multiple sites may be selected for replication under a statistical significance cutoff rule. In simulations calibrated to the JOBSTART demonstration, we find that multiple sites can be selected as statistically significant with a probability of between $17.6\%$ and $53.4\%$, depending on our choice of significance level by which to select sites for replication.
@@ 3.1 JOBSTART: Empirical Findings
The JOBSTART demonstration compared 13 distinct implementations of a job training and retraining program targeted towards low-skilled school dropouts at different sites. Out of these 13 interventions, only one intervention, the CET at San Jose, had a statistically significant effect on earnings, thus motivating the replication in Miller et al. (2005). In the table below, we report the effects of the JOBSTART intervention on earnings at each of these 13 sites, with point estimates due to Cave et al. (1993) and standard errors due to Andrews et al. (2023):
@@
Program treatment effects from the JOBSTART demonstration, as reported in Cave et al. (1993). The reported standard errors are those derived in Andrews et al. (2023).
In what follows, we denote the ATEs for the thirteen interventions by $Y$, and index $Y$ by $j$ in the set of interventions $J$. Of all the interventions from the JOBSTART demonstration, only the CET program had a statistically significant treatment effect on earnings at the $1\%$ level. [^2] Or, similarly, at the $5\%$ level. In the thought experiment where selection for replication is based on statistical significance, the conditional inference procedures from Andrews et al. (2023) must be revised. However, the econometrician does not know, ex-ante, how many studies will be selected for replication as statistically significant. Consequently, the econometrician will choose to apply conditional inference only when one study is statistically significant. As a result, the conditioning event chosen for conditional inference must account for the fact that the econometrician’s choice of inference procedure is itself data-dependent. That is, we condition on the event that only the winner $\jmath$ in $J$ is statistically significant at a given level. Indeed:
$Y_{\jmath}/\sqrt{\Sigma_{Y,\jmath\jmath}}\geq c,\quad Y_{j}/\sqrt{\Sigma_{Y,jj}}\leq c\quad\text{for all }j\neq\jmath$ (4)
We can construct a confidence set conditionally valid for the event 4 using the approaches developed in Andrews et al. (2023), since their propositions 1-3 and their proposition 7 can easily be applied to accomodate selection events of the form above. We present our empirical findings using our novel two-step method, the projection approach, and both the original and revised conditional and hybrid approaches. We find that the two-step approach provides confidence sets that are between 5% and 8% narrower than the projection approach. A back of the envelope calculation gives that our two-step confidence region increases the lower bound on total effects of the replication of Miller et al. (2005) by approximately \$122,000. The revised conditional and hybrid approaches provide narrower-still confidence sets, but may be of limited use in the JOBSTART application due to the frequency with which we may observe statistically significant outcomes for multiple interventions.
@@
Confidence sets for the CET program, correcting for cutoff-based selection. Confidence sets are presented for the 5% significance cutoff.
@@
We now present results from simulations calibrated to the JOBSTART demonstration. We demonstrate that the probability of making multiple selections when using a cutoff rule is non-negligible. In such instances, conditional and hybrid confidence sets do not readily apply, as discussed in 4. In the simulation presented below, we compute the probability of making multiple selections for 1000 simulation draws. We find that in 90% of our simulations, when choosing a 1% significance cutoff rule for selection, the probability of making multiple selections lies between 0.153 and 0.828. When choosing a 5% significance cutoff, this range becomes the interval $[0.473,0.977]$. We present a histogram of these probabilities, over a confidence region for the means $\mu$ given our observed data, below:

### v4

@@ 5 Application: the JOBSTART Demonstration
In this section, we revisit the JOBSTART demonstration due to Cave et al. (1993) and subsequent replication failure in Miller et al. (2005), which has been previously studied in the selective inference literature by Andrews et al. (2023). The JOBSTART demonstration was a randomized controlled trial taking place between 1985 and 1988 across 13 sites, with the intention of studying the effects of a vocational training program on the employment outcomes of young, low-skilled school dropouts. The treatment group was given access to a suite of JOBSTART services which were inaccessible to those in the control group. Among the sites included in the JOBSTART study, only one site, the Center for Employment Training (CET) in San Jose, saw a large and statistically significant estimate of the effect on earnings. Cave et al. (1993) note that they cannot attribute the unique success of CET to a particular feature of the program, but suggest that the CET’s strong connections with San Jose employers, or their robust placement efforts, may explain some of its value-add. Motivated by the success of the CET program, Miller et al. (2005) replicate the intervention at 12 sites. They find that, even in replication sites deemed to have high fidelity to the original CET program of Cave et al. (1993), the estimated effect of the program’s services on enrollees’ earnings was not statistically significant.
@@
Andrews et al. (2023) study the possibility that this replication failure is due to a winner’s curse. In particular, they consider the possibility that the estimates of the effect of CET on earnings in the original JOBSTART demonstration are upwardly biased by virtue of CET being the site with the largest estimated effect. We consider a complementary thought experiment. In particular, since the exact mechanism by which Miller et al. (2005) select a program for replication is unknown, we consider the possibility that the replication failure between the two studies can be explained by an alternative selection rule in which the sites selected by Miller et al. (2005) are chosen according to a statistical significance cutoff. Under this selection rule, it is impossible for the analyst to know ex-ante how many sites will be selected for replication. When multiple selections are made, the conditional and hybrid approaches of Andrews et al. (2023) are difficult to apply, as we discuss in section 4.1.555It is not difficult to provide conditional and hybrid confidence sets that apply in the event that only one selection is made via a particular choice of conditioning set. Such concerns are discussed in appendix C of Andrews et al. (2023), but are beyond the scope of our analysis here.
@@
Empirically, we show that our two-step confidence regions for the effect of the CET program on earnings exclude zero. This finding suggests that a winner’s curse does not fully explain the replication failure between the studies of Cave et al. (1993) and Miller et al. (2005), even under an alternative selection rule. Finally, motivated by concerns that conditional inference may be most appropriate when one selection is made, we illustrate the frequency with which multiple sites may be selected for replication under a statistical significance cutoff rule. In simulations calibrated to the JOBSTART demonstration, we find that multiple sites can be selected as statistically significant with a mean probability of between $17.6\%$ and $53.4\%$, depending on our choice of significance level by which to select sites for replication.
@@ 5.1 JOBSTART: Empirical Findings
In the table below, we report the effects of the JOBSTART intervention on earnings at each of these 13 sites, with point estimates due to Cave et al. (1993) and standard errors due to Andrews et al. (2023):
@@
Estimated program treatment effects from the JOBSTART demonstration, as reported in Cave et al. (1993). The reported standard errors are those derived in Andrews et al. (2023).
In what follows, we denote the estimated ATEs for the thirteen interventions by $Y$, and index $Y$ by $j$ in the set of interventions $J$. Of all the interventions from the JOBSTART demonstration, only the CET program had a statistically significant treatment effect on earnings at the $1\%$ level. [^6] Or, similarly, at the $5\%$ level. In the thought experiment where Miller et al. (2005) select statistically significant programs for replication rather than the winner, we provide confidence sets for the true treatment effects at selected sites.
We present our empirical findings using both the symmetric and asymmetric versions of our novel two-step method. We also provide a conditional confidence set under the winner selection rule from Andrews et al. (2023). We emphasize, however, that the winner selection rule leads to a different parameter and notion of coverage, so conditional inference is not directly comparable in this setting.
@@
Confidence sets for the CET program, correcting for cutoff-based selection. Confidence sets are presented for a 5% significance cutoff. The conditional confidence set presented assumes the original, best-treatment selection rule considered in Andrews et al. (2023).
@@
We now present results from simulations calibrated to the JOBSTART demonstration. We demonstrate that the probability of making multiple selections when using a cutoff rule is non-negligible. In such instances, conditional and hybrid confidence sets do not readily apply, as explained in section 4. In the simulation presented below, we compute the probability of making multiple selections for 1000 simulation draws. We find that in 90% of our simulations, when choosing a 1% significance cutoff rule for selection, the probability of making multiple selections lies between 0.153 and 0.828. When choosing a 5% significance cutoff, this range becomes the interval $[0.473,0.977]$. We present a histogram of these probabilities, over a confidence region for the means $\mu$ given our observed data, below: